12 GeV Upgrade
At the Hall D construction site, excavation for the tunnel connection is now approximately 10 feet deep, which is two feet above the existing tunnel. The excavation is an iterative process with the dewatering system. As the excavation goes deeper, additional dewatering wells are installed to further lower the water table in the area of the excavation. Next, the contractor will be installing a type of shoring called a trench box at the existing northeast stub to establish safe working conditions in the excavation for the new tunnel extension. The contractor is also installing the metal siding and roofing on the Service Building, which will house the power supplies for the Hall D beamline and tagger magnets, instrumentation racks and mechanical cooling systems.
Hall C staffers have started work on removing parts of the Short Orbit Spectrometer in preparation for its full removal next year. Signal cable, DC power and water hoses are being removed from cable trays and the flexible cable tubes under the spectrometer. After removal, the signal cables will be evaluated for possible reuse in the future Super High Momentum Spectrometer.
The summer scheduled accelerator down went into the second week, with tasks performed as planned. The preparation for the Polarized Electrons for Polarized Positrons project started. Removal and installation of west arc dipoles continued. Microphonics and vibration measurements were conducted on the newly installed R100 cryomodule at 2L21. Many other tasks were also performed. On Friday, a no-prior-notice emergency response exercise on the accelerator site was conducted; Newport News Fire and Rescue units participated in this exercise.
The FEL got the new low-conductivity cooling water system turned on again two weeks early; after it circulates for cleanup, the FEL team can put it through the FEL again. The team is engaging with vacuum and superconducting radiofrequency groups to get the FEL gun changed out; this effort is complicated by two leaky beamline valves that could contaminate the booster and also helium leaks in the booster, which makes the group nervous about warming the system up for fear of rendering it totally inoperable.
The cryogenics group has been working with the Hall D solenoid magnet. The third coil of the four coil set has successfully passed the testing program. The coil was ramped to 1500 amps several times without incident. One coil remains to be tested.
Facilities Management and Logistics
All domestic water in the ARC will be turned off for a one-hour period beginning at 6 p.m. today, while a water system leak is repaired. During this time, there will be no water in the restrooms, kitchenettes, water fountains, labs, etc.
JLab Calendar of Events
JLab Summer 2011 Education Calendar
May 31- June 17: Hampton University Graduate Studies
Environment, Safety, Health and Quality
Electrical Cord Condition – at Work and at Home
Computing and Networking Infrastructure
External DNS Server Outage
ESnet Upgrade Set for June 7
For more information, including the latest status and contact information, visit the Library webpage.
Fire Academy Graduation Ceremony
Sign Up Your Kids for a Summer Physics Fest
Tracking Thomas at JLab
Congratulations to David Fazenbaker, who was the first person to correctly identify the May 25 location. Honorable mentions go out to Jason Willoughby, Dena Polyhronakis, Chandra Gilchrist, Stephanie Vermeire, Sharon Hay and Connie Adams. Check out the Tracking Thomas webpage for a better view of his last location and this week's new mystery photo.